OpenStudy (anonymous):

\[\left(\begin{matrix}9x \\x^2-4\end{matrix}\right)+\left(\begin{matrix}x \\ x-2\end{matrix}\right)\] Add and simplify

myininaya (myininaya):

ok so it looks like you are working with a matrix

myininaya (myininaya):

oh wow it didn't show up until now thats crazy

OpenStudy (anonymous):

lol sorry

myininaya (myininaya):

just add straight across

myininaya (myininaya):

\[\left[\begin{matrix}(9x)+(x) & \\ (x^2-4)+(x+2)&\end{matrix}\right]\]

myininaya (myininaya):

\[=\left[\begin{matrix}10x & \\ x^2+x-2 &\end{matrix}\right]\]

OpenStudy (anonymous):

ok so would it be \[\left(\begin{matrix}10x ^{} \\ 2x ^{2}-2\end{matrix}\right)\]